# Error Bound Analysis for the Regularized Loss of Deep Linear Neural Networks

### v1

@@ 4.2 Proofs of the Error Bound
$\geq\lambda\sigma_{r_{\sigma}+1}(\bm{W}_{l})-\sqrt{\lambda}y_{1}\sigma_{r_{\sigma}+1}^{L-1}(\bm{W}_{l})\geq\sqrt{\lambda}\left(\sqrt{\lambda}-y_{1}\sigma_{r_{\sigma}+1}^{L-2}(\bm{W}_{l})\right)\sigma_{r_{\sigma}+1}(\bm{W}_{l}),$ (151)
@@
where $\bm{\Psi}_{r_{\sigma}+1:d_{L},r_{\sigma}+1:d_{0}}$ denotes the submatrix of $\bm{\Psi}$ with the rows being indexed from $r_{\sigma}+1$ to $d_{L}$ and the columns being indexed from $r_{\sigma}+1$ to $d_{0}$, and the third inequality is due to $\|\bm{\Psi}_{r_{\sigma}+1:d_{L},r_{\sigma}+1:d_{0}}\|\leq\|\bm{Y}\|=y_{1}$. Using Weyl’s inequality and (79) with $L\geq 3$, we have
@@
$\sigma_{r_{\sigma}+1}(\bm{W}_{l})\leq\mathrm{dist}\left(\bm{W},\mathcal{W}_{\bm{\sigma}^{*}}\right)\leq\left(\frac{\sqrt{\lambda}}{2y_{1}}\right)^{1/(L-2)}.$
This, together with (4.2), directly yields $\sigma_{r_{\sigma}+1}(\bm{W}_{l})\leq{2\eta_{2}}\|\nabla G(\bm{W})\|_{F}/{\lambda}$, which further implies (80). ∎
Equipped with Lemma 3, 2, 4, Lemma 4, 5, and Davis-Kahan Theorem (see Lemma 10), we are ready to prove 6.
Proof of 6.
According to (77a) or (79), we have $\mathrm{dist}(\bm{W},\mathcal{W})\leq\delta_{\sigma}/3<\sigma_{\min}^{*}/2$. This, together with Lemma 3 and 2, implies that (57)-(60) and (63) hold. According to 4 and Lemma 4 with $\mathrm{dist}(\bm{W},\mathcal{W})\leq\delta_{\sigma}/3$, there exist matrices $\bm{T}_{l}^{(i)}\in\mathcal{O}^{g_{i}}$ for each $i\in[p]$ and $\bm{T}_{l}^{(p+1)}\in\mathcal{O}^{d_{l-1}-r_{\sigma}}$ such that (69) and (70) hold, where $\bm{A}_{i}$ and $\bm{B}_{i}$ for each $i\in[p+1]$ are defined in (71) and (72). According to (77) (resp., (79)) and 5, we have (78) (resp., (80)) hold. Recall that (66) denotes an SVD of $\bm{W}_{l}$ for each $l\in[L]$. For ease of exposition, let
$\bm{\Psi}:=\bm{U}_{L}^{T}\bm{Y}\bm{V}_{1},\ \hat{\bm{\Sigma}}_{l}:=\mathrm{BlkDiag}\left(\bm{\Sigma}_{l}^{(1)},\dots,\bm{\Sigma}_{l}^{(p)},\bm{0}_{(d_{l}-r_{\sigma})\times(d_{l-1}-r_{\sigma})}\right),\ \forall l\in[L].$ (152)
where $\bm{\Sigma}_{l}^{(1)},\dots,\bm{\Sigma}_{l}^{(p)}$ are defined in (66) for each $l\in[L]$. Using the form of $\bm{\Sigma}_{l}$ in (66) and (78) (resp., (80)), we have for each $l\in[L]$,
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\quad\ \leq c_{3}\sqrt{d_{\max}}\left(\frac{3\sigma_{\max}^{*}}{2}\right)^{2(L-1)}\|\nabla G(\bm{W})\|_{F},$ (154)
where the inequality uses (57) and (153).
(i) We first focus on $l=L$. Now, we bound
$\left\|(\hat{\bm{\Sigma}}_{L}\hat{\bm{\Sigma}}_{L}^{T})^{L-1}\hat{\bm{\Sigma}}_{L}+\lambda\hat{\bm{\Sigma}}_{L}-\sqrt{\lambda}\bm{\Psi}\mathrm{BlkDiag}\left(\bm{B}_{1},\dots,\bm{B}_{p},\bm{0}\right)\right\|_{F}$
$\leq$ $\left\|(\bm{\Sigma}_{L}\bm{\Sigma}_{L}^{T})^{L-1}\bm{\Sigma}_{L}+\lambda\bm{\Sigma}_{L}-\sqrt{\lambda}\bm{\Psi}\mathrm{BlkDiag}\left(\bm{B}_{1},\dots,\bm{B}_{p},\bm{B}_{p+1}\right)\right\|_{F}+\lambda\|\hat{\bm{\Sigma}}_{L}-\bm{\Sigma}_{L}\|_{F}$
$\leq$ $\left(c_{2}+\left(\left(\frac{3\sigma_{\max}^{*}}{2}\right)^{2(L-1)}+\lambda\right)c_{3}\sqrt{d_{\max}}\right)\left\|\nabla G(\bm{W})\right\|_{F}+\sqrt{\lambda}y_{1}\|\bm{B}_{p+1}\|_{F}\leq\eta_{3}\left\|\nabla G(\bm{W})\right\|_{F},$ (155)
where the second inequality follows from (70), (72), (87), (153), and (154), and the last inequality uses (57), (72), (153), and
$\eta_{3}:=c_{2}+c_{3}\sqrt{d_{\max}}\left(\left(\frac{3\sigma_{\max}^{*}}{2}\right)^{2(L-1)}+\sqrt{\lambda}y_{1}\left(\frac{3\sigma_{\max}^{*}}{2}\right)^{L-2}+\lambda\right).$ (156)
Substituting $\hat{\bm{\Sigma}}_{L}$ in (160), (72), and (87) into (4.2) yields
$\leq$ $\left\|(\hat{\bm{\Sigma}}_{L}\hat{\bm{\Sigma}}_{L}^{T})^{L-1}\hat{\bm{\Sigma}}_{L}+\lambda\hat{\bm{\Sigma}}_{L}-\sqrt{\lambda}\bm{\Psi}\hat{\bm{T}}\mathrm{BlkDiag}\left((\bm{\Sigma}_{L}^{(1)})^{L-1},\dots,(\bm{\Sigma}_{L}^{(p)})^{L-1},\bm{0}\right)\right\|_{F}\leq\eta_{3}^{2}\|\nabla G(\bm{W})\|_{F}^{2}.$ (157)
Using this and the definition of $\varphi(\cdot)$ in (82), together with dividing the above inequality by $\sqrt{\lambda}\sigma_{r_{\sigma}}^{L-1}(\bm{W}_{L})$ on both sides, yields
$\sum_{i=1}^{r_{\sigma}}\left\|\varphi(\sigma_{i}(\bm{W}_{L}))\bm{e}_{i}-\bm{\Psi}\hat{\bm{T}}\bm{e}_{i}\right\|^{2}$ $\leq\frac{\eta_{3}^{2}}{\lambda\sigma_{r_{\sigma}}^{2(L-1)}(\bm{W}_{L})}\|\nabla G(\bm{W})\|_{F}^{2}\leq\frac{2^{2(L-1)}\eta_{3}^{2}}{\lambda(\sigma_{\min}^{*})^{2L-2}}\|\nabla G(\bm{W})\|_{F}^{2},$ (158)
where the last inequality uses (58). Using the same argument of (4.2) for (69), we have
$\left\|(\hat{\bm{\Sigma}}_{1}\hat{\bm{\Sigma}}_{1}^{T})^{L-1}\hat{\bm{\Sigma}}_{1}+\lambda\hat{\bm{\Sigma}}_{1}-\sqrt{\lambda}\mathrm{BlkDiag}\left(\bm{A}_{1},\dots,\bm{A}_{p},\bm{0}\right)\bm{\Psi}\right\|_{F}\leq\eta_{3}\left\|\nabla G(\bm{W})\right\|_{F}.$ (159)
Then we compute
$\left\|\left(\hat{\bm{\Sigma}}_{1}\hat{\bm{\Sigma}}_{1}^{T}\right)^{L-1}\hat{\bm{\Sigma}}_{1}+\lambda\hat{\bm{\Sigma}}_{1}-\sqrt{\lambda}\,\mathrm{BlkDiag}\left((\bm{\Sigma}_{1}^{(1)})^{L-1},\dots,(\bm{\Sigma}_{1}^{(p)})^{L-1},\bm{0}\right)\bm{\Psi}\hat{\bm{T}}\right\|_{F}\leq\eta_{4}\left\|\nabla G(\bm{W})\right\|_{F},$ (160)
@@
$\eta_{4}:=\eta_{3}+\frac{p\eta_{1}(2L-1)L}{\sigma^{*}_{\min}}\left(\frac{3\sigma^{*}_{\max}}{2}\right)^{2L-2}+\frac{\lambda p\eta_{1}L}{\sigma^{*}_{\min}}.$
To maintain the flow of the main proof, we defer the proof of (160) to Appendix D.1. Applying a similar argument as in (158) to rows of the above inequality yields
$\sum_{i=1}^{r_{\sigma}}\left\|\varphi(\sigma_{i}(\bm{W}_{1}))\bm{e}_{i}^{T}-\bm{e}_{i}^{T}\bm{\Psi}\hat{\bm{T}}\right\|^{2}$ $\leq\left(\frac{2^{L-1}\eta_{4}}{\sqrt{\lambda}\sigma_{\min}^{*{L-1}}}\right)^{2}\|\nabla G(\bm{W})\|_{F}^{2}.$ (161)
To proceed, we define $\bm{d}:=\left(\varphi(\sigma_{1}(\bm{W}_{L})),\dots,\varphi(\sigma_{r_{\sigma}}(\bm{W}_{L}))\right)\in\mathbb{R}^{r_{\sigma}}$ and
$\bm{Z}:=\mathrm{BlkDiag}\left(\mathrm{diag}(\bm{d}),(\bm{\Psi}\hat{\bm{T}})_{r_{\sigma}+1:d_{L},r_{\sigma}+1:d_{0}}\right),$
where $(\bm{\Psi}\hat{\bm{T}})_{r_{\sigma}+1:d_{L},r_{\sigma}+1:d_{0}}$ denotes a submatrix of $\bm{\Psi}\hat{\bm{T}}$ with rows indexed by $r_{\sigma}+1:d_{L}$ and columns indexed by $r_{\sigma}+1:d_{0}$. Now, we compute
@@
$=\sum_{i=1}^{r_{\sigma}}\left\|\bm{\Psi}\hat{\bm{T}}\bm{e}_{i}-\varphi(\sigma_{i}(\bm{W}_{L}))\bm{e}_{i}\right\|^{2}+\left\|\left(\bm{\Psi}\hat{\bm{T}}\right)_{1:r_{\sigma},r_{\sigma}+1:d_{0},}\right\|_{F}^{2}$
@@
$\leq\left(\frac{2^{L-1}}{\sqrt{\lambda}\sigma_{\min}^{*{L-1}}}\right)^{2}\left(\eta_{3}^{2}+\eta_{4}^{2}\right)\|\nabla G(\bm{W})\|_{F}^{2},$ (162)
where the second equality uses the structure of $\bm{Z}$, the first inequality is due to
$\left\|\bm{e}_{i}^{T}\bm{\Psi}\hat{\bm{T}}-\varphi(\sigma_{i}(\bm{W}_{1}))\bm{e}_{i}^{T}\right\|^{2}\geq\left\|\left(\bm{\Psi}\hat{\bm{T}}\right)_{i,r_{\sigma}+1:d_{0}}\right\|^{2},\ \forall i\in[r_{\sigma}],$
@@
and the second inequality follows from (158) and (161). Let $(\bm{\Psi}\hat{\bm{T}})_{r_{\sigma}+1:d_{L},r_{\sigma}+1:d_{0}}=\bm{P}\bm{\Lambda}\bm{Q}^{T}$ be an SVD of $(\bm{\Psi}\hat{\bm{T}})_{r_{\sigma}+1:d_{L},r_{\sigma}+1:d_{0}}$, where $\bm{P}\in\mathcal{O}^{d_{L}-r_{\sigma}}$, $\bm{Q}\in\mathcal{O}^{d_{0}-r_{\sigma}}$, and $\bm{\Lambda}\in\mathbb{R}^{(d_{L}-r_{\sigma})\times(d_{0}-r_{\sigma})}$ with diagonal elements $\gamma_{1}\geq\dots\geq\gamma_{d_{\min}-r_{\sigma}}$ being top $d_{\min}-r_{\sigma}$ singular values. Now, we define
@@
$\bm{c}:=\left(\varphi(\sigma_{1}(\bm{W}_{L})),\dots,\varphi(\sigma_{r_{\sigma}}(\bm{W}_{L})),\gamma_{1},\dots,\gamma_{d_{\min}-r_{\sigma}}\right)\in\mathbb{R}^{d_{\min}}.$ (163)
@@
Let $\bm{\Pi}\in\mathcal{P}^{d_{\min}}$ be a permutation matrix such that the entries $\bm{\Pi}\bm{c}$ are in a decreasing order and
$\bar{\bm{U}}_{L}:=\bm{U}_{L}\mathrm{BlkDiag}\left(\bm{I}_{r_{\sigma}},\bm{P}^{T}\right)\mathrm{BlkDiag}\left(\bm{\Pi}^{T},\bm{I}_{d_{L}-d_{\min}}\right),$
$\bar{\bm{V}}_{1}:=\bm{V}_{1}\hat{\bm{T}}\mathrm{BlkDiag}\left(\bm{I}_{r_{\sigma}},\bm{Q}^{T}\right)\mathrm{BlkDiag}\left(\bm{\Pi}^{T},\bm{I}_{d_{0}-d_{\min}}\right),$ (164)
$\bar{\bm{Y}}:=\mathrm{BlkDiag}\left(\bm{\Pi}\mathrm{diag}\left(\bm{c}\right)\bm{\Pi}^{T},\bm{0}_{(d_{L}-d_{\min})\times(d_{0}-d_{\min})}\right).$
@@
Since both $\bm{Y}$ and $\bar{\bm{Y}}$ are diagonal matrices with elements in decreasing order, by using Mirsky’s inequality (see Lemma 11), we have
$\|\bm{Y}-\bar{\bm{Y}}\|_{F}\leq\|\bar{\bm{U}}_{L}^{T}\bm{Y}\bar{\bm{V}}_{1}-\bar{\bm{Y}}\|_{F}\leq\frac{2^{L-1}}{\sqrt{\lambda}\sigma_{\min}^{*{L-1}}}\sqrt{\eta_{3}^{2}+\eta_{4}^{2}}\|\nabla G(\bm{W})\|_{F},$ (165)
@@
where the last inequality uses (4.2). This, together with (83b), implies
$\|\bm{Y}-\bar{\bm{Y}}\|_{F}\leq\frac{\delta_{y}}{3},\ \|\bar{\bm{Y}}\|\leq\|\bar{\bm{Y}}-\bm{Y}\|+\|\bm{Y}\|\leq y_{1}+\frac{\delta_{y}}{3}.$ (166)
According to the triangular inequality, we obtain
$\|\bar{\bm{U}}_{L}^{T}\bm{Y}\bm{Y}^{T}\bar{\bm{U}}_{L}-\bar{\bm{Y}}\bar{\bm{Y}}^{T}\|_{F}$ $\leq\|(\bar{\bm{U}}_{L}^{T}\bm{Y}\bar{\bm{V}}_{1}-\bar{\bm{Y}})\bar{\bm{V}_{1}}^{T}\bm{Y}^{T}\bar{\bm{U}}_{L}\|_{F}+\|\bar{\bm{Y}}(\bar{\bm{U}}_{L}^{T}\bm{Y}\bar{\bm{V}}_{1}-\bar{\bm{Y}})^{T}\|_{F}$

### v3

@@ 3.2.2 Proof of the Error Bound
$\geq$ $\ \lambda\sigma_{r_{\sigma}+1}(\bm{W}_{l})-\sqrt{\lambda}y_{1}\sigma_{r_{\sigma}+1}^{L-1}(\bm{W}_{l})\geq\sqrt{\lambda}\left(\sqrt{\lambda}-y_{1}\sigma_{r_{\sigma}+1}^{L-2}(\bm{W}_{l})\right)\sigma_{r_{\sigma}+1}(\bm{W}_{l}),$
@@
where $\bm{\Psi}_{r_{\sigma}+1:d_{L},r_{\sigma}+1:d_{0}}$ denotes the submatrix of $\bm{\Psi}$ with the rows indexed from $r_{\sigma}+1$ to $d_{L}$ and the columns indexed from $r_{\sigma}+1$ to $d_{0}$, and the third inequality is due to $\|\bm{\Psi}_{r_{\sigma}+1:d_{L},r_{\sigma}+1:d_{0}}\|\leq\|\bm{Y}\|=y_{1}$. Using Weyl’s inequality and (72), we have $\sigma_{r_{\sigma}+1}(\bm{W}_{l})\leq\mathrm{dist}\left(\bm{W},\mathcal{W}_{\bm{\sigma}^{*}}\right)\leq\left(\frac{\sqrt{\lambda}}{2y_{1}}\right)^{1/(L-2)}.$ This, together with the above inequality , directly yields $\sigma_{r_{\sigma}+1}(\bm{W}_{l})\leq{2\eta_{2}}\|\nabla G(\bm{W})\|_{F}/{\lambda}$, which implies (73). ∎
The above proposition provides bounds for the $\min\{d_{l},d_{l-1}\}-r_{\sigma}$ smallest singular values of $\bm{W}_{l}$. It remains to bound the top $r_{\sigma}$ singular values and all singular vectors. For convenience, we define an auxiliary function $\varphi:\mathbb{R}_{++}\;\to\;\mathbb{R}$:
$\varphi(x):=\frac{x^{2L-1}+\lambda x}{\sqrt{\lambda}\,x^{L-1}}.$ (85)
Proposition 5.
Let $\bm{\sigma}^{*}\in\mathcal{A}_{\rm sort}\setminus\{\bm{0}\}$ be arbitrary and $\bm{W}$ be arbitrary such that
@@
$\mathrm{dist}(\bm{W},\mathcal{W}_{\bm{\sigma}^{*}})\leq\min\left\{\delta_{1},\delta_{2},\frac{\delta_{\sigma}}{3}\right\}\quad\text{and}\quad\|\nabla G(\bm{W})\|_{F}\leq\frac{\delta_{y}\sqrt{\lambda}(\sigma^{*}_{\min})^{L-1}}{3\cdot 2^{L-1}\sqrt{\eta_{3}^{2}+\eta_{4}^{2}}},$ (86)
@@
$\delta_{1}$ $:=\frac{\delta_{y}}{{3L}\left(\frac{4\sigma_{\max}^{*}}{3}\right)^{L-1}\frac{1}{\sqrt{\lambda}}+3(L-2)\sqrt{\lambda}\left(\frac{2\sigma_{\min}^{*}}{3}\right)^{1-L}},$
$\delta_{2}$ $:=\frac{\min_{i\in[r_{\sigma}]}|\varphi^{\prime}(\sigma_{i}^{*})|}{{2L(L-1)}\left(\frac{4\sigma_{\max}^{*}}{3}\right)^{L-1}\frac{1}{\sqrt{\lambda}}+2\sqrt{\lambda}(2-L)(1-L)\left(\frac{2\sigma_{\min}^{*}}{3}\right)^{-L}},$ (87)
$\ \eta_{3}$ $:=c_{2}+c_{3}\sqrt{d_{\max}}\left(\left(\frac{3\sigma_{\max}^{*}}{2}\right)^{2(L-1)}\!\!\!\!+\sqrt{\lambda}y_{1}\left(\frac{3\sigma_{\max}^{*}}{2}\right)^{L-2}\!\!\!\!+\lambda\right),$
$\eta_{4}$ $:=\eta_{3}+\frac{p\eta_{1}(2L-1)L}{\sigma^{*}_{\min}}\left(\frac{3\sigma^{*}_{\max}}{2}\right)^{2L-2}\!\!\!\!+\frac{\lambda p\eta_{1}L}{\sigma^{*}_{\min}}.$ (88)
Suppose in addition Assumptions 1 and 2 hold, (70) holds for $L=2$, and (72) holds for $L\geq 3$.
(i) There exist orthogonal matrices $\hat{\bm{U}}_{L}^{(i)}\in\mathcal{O}^{h_{i}}$ for $i\in[p_{Y}]$, $\hat{\bm{U}}_{L}^{(p_{Y}+1)}\in\mathcal{O}^{d_{L}-r_{Y}}$, $\hat{\bm{V}}_{L}^{(p_{Y}+1)}\in\mathcal{O}^{d_{0}-r_{Y}}$, $\bm{T}_{l}^{(i)}\in\mathcal{O}^{g_{i}}$ for $i\in[p]$, $\bm{P}\in\mathcal{O}^{d_{L}-r_{\sigma}}$, $\bm{Q}\in\mathcal{O}^{d_{0}-r_{\sigma}}$, and a permutation matrix $\bm{\Pi}\in\mathcal{P}^{d_{\min}}$ with $(\bm{\sigma}^{*},\bm{\Pi}^{T})\in\mathcal{B}$ such that
$\|\tilde{\bm{U}}_{L}-\bm{U}_{L}\|_{F}\leq c_{4}\|\nabla G(\bm{W})\|_{F}\quad\text{ and }\quad\|\tilde{\bm{V}}_{1}-\bm{V}_{1}\|_{F}\leq c_{4}\|\nabla G(\bm{W})\|_{F},$ (89)
where
$\eta_{5}:=\frac{2^{L+1}(6y_{1}+\delta_{y})(p_{Y}+1)\sqrt{\eta_{3}^{2}+\eta_{4}^{2}}}{3\sqrt{\lambda}y_{s_{p_{Y}}}\delta_{y}\sigma^{*L-1}_{\min}},\ c_{4}:=\eta_{5}+\frac{1}{y_{s_{p_{Y}}}}\left(\frac{2^{L}\sqrt{\eta_{3}^{2}+\eta_{4}^{2}}}{\sqrt{\lambda}\sigma^{*L-1}_{\min}}+2y_{1}\eta_{5}\right)$ (90)
$\tilde{\bm{U}}_{L}:=\mathrm{BlkD}\left(\hat{\bm{U}}_{L}^{(1)},\dots,\hat{\bm{U}}_{L}^{(p_{Y})},\hat{\bm{U}}_{L}^{(p_{Y}+1)}\right)\mathrm{BlkD}\left(\bm{\Pi},\bm{I}_{d_{L}-d_{\min}}\right)\mathrm{BlkD}\left(\bm{I}_{r_{\sigma}},\bm{P}^{T}\right),$ (91)
$\tilde{\bm{V}}_{1}:=\mathrm{BlkD}\left(\hat{\bm{U}}_{L}^{(1)},\dots,\hat{\bm{U}}_{L}^{(p_{Y})},\hat{\bm{V}}_{L}^{(p_{Y}+1)}\right)\mathrm{BlkD}\left(\bm{\Pi},\bm{I}_{d_{0}-d_{\min}}\right)\mathrm{BlkD}\left(\bm{I}_{r_{\sigma}},\bm{Q}^{T}\right)\hat{\bm{T}}$, (92)
(ii) For each $l\in[L]$, there exists a constant $c_{5}>0$ such that
$|\sigma_{i}\left(\bm{W}_{l})-\sigma^{*}_{i}\right|\leq c_{5}\|\nabla G(\bm{W})\|_{F},\ \forall i\in[r_{\sigma}],$ (94)
where
$c_{5}:=\frac{2^{L}\sqrt{\eta_{3}^{2}+\eta_{4}^{2}}}{\sqrt{\lambda}\sigma_{\min}^{*{L-1}}\min_{i\in[r_{\sigma}]}\left|\varphi^{\prime}(\sigma_{i}^{*})\right|}+\frac{3\sqrt{2}(L-1)\sigma_{\max}^{*}}{4\lambda\sigma_{\min}^{*}}.$ (95)
Proof.
According to (70a) and (72), we have $\mathrm{dist}(\bm{W},\mathcal{W}_{\bm{\sigma}^{*}})\leq\delta_{\sigma}/3<\sigma_{\min}^{*}/2$. This, together with Lemma 2 and 2, implies that (35)-(37) and (41) hold. According to 3 and Lemma 3 with $\mathrm{dist}(\bm{W},\mathcal{W}_{\bm{\sigma}^{*}})\leq\delta_{\sigma}/3$, there exist matrices $\bm{T}_{l}^{(i)}\in\mathcal{O}^{g_{i}}$ for each $i\in[p]$ and $\bm{T}_{l}^{(p+1)}\in\mathcal{O}^{d_{l-1}-r_{\sigma}}$ such that (58) and (59) hold, where $\bm{A}_{i}$ and $\bm{B}_{i}$ for $i\in[p+1]$ are defined in (54) and (55). Based on (70) (resp., (72)) and 4, we have (71) (resp., (73)). Recall that (3.2.1) denotes an SVD of $\bm{W}_{l}$ for $l\in[L]$. Let
$\bm{\Psi}:=\bm{U}_{L}^{T}\bm{Y}\bm{V}_{1},\qquad\hat{\bm{\Sigma}}_{l}:=\mathrm{BlkD}\left(\bm{\Sigma}_{l}^{(1)},\dots,\bm{\Sigma}_{l}^{(p)},\bm{0}_{(d_{l}-r_{\sigma})\times(d_{l-1}-r_{\sigma})}\right),\ \forall l\in[L].$ (96)
(i) We first prove the first inequality in (89). We claim that the following two inequalities hold:
$\sum_{i=1}^{r_{\sigma}}\left\|\varphi(\sigma_{i}(\bm{W}_{L}))\bm{e}_{i}-\bm{\Psi}\hat{\bm{T}}\bm{e}_{i}\right\|^{2}$ $\leq\frac{4^{L-1}\eta_{3}^{2}}{\lambda(\sigma_{\min}^{*})^{2L-2}}\|\nabla G(\bm{W})\|_{F}^{2},$ (97)
$\sum_{i=1}^{r_{\sigma}}\left\|\varphi(\sigma_{i}(\bm{W}_{1}))\bm{e}_{i}^{T}-\bm{e}_{i}^{T}\bm{\Psi}\hat{\bm{T}}\right\|^{2}$ $\leq\frac{4^{L-1}\eta_{4}^{2}}{\lambda(\sigma_{\min}^{*})^{{2L-2}}}\|\nabla G(\bm{W})\|_{F}^{2}.$ (98)
To maintain the flow of the main proof, the proof of this claim is deferred to Appendix B.3.1. Now, let $\bm{\alpha}:=\left(\varphi(\sigma_{1}(\bm{W}_{L})),\dots,\varphi(\sigma_{r_{\sigma}}(\bm{W}_{L}))\right)\in\mathbb{R}^{r_{\sigma}}$ and $\bm{Z}:=\mathrm{BlkDiag}\left(\mathrm{diag}(\bm{\alpha}),(\bm{\Psi}\hat{\bm{T}})_{r_{\sigma}+1:d_{L},r_{\sigma}+1:d_{0}}\right)$. We have
@@
$=\sum_{i=1}^{r_{\sigma}}\left\|\bm{\Psi}\hat{\bm{T}}\bm{e}_{i}-\varphi(\sigma_{i}(\bm{W}_{L}))\bm{e}_{i}\right\|^{2}+\left\|\left(\bm{\Psi}\hat{\bm{T}}\right)_{1:r_{\sigma},r_{\sigma}+1:d_{0}}\right\|_{F}^{2}$
@@
[…]
where the second equality uses the structure of $\bm{Z}$ and the first inequality is due to
$\left\|\bm{e}_{i}^{T}\bm{\Psi}\hat{\bm{T}}-\varphi(\sigma_{i}(\bm{W}_{1}))\bm{e}_{i}^{T}\right\|\geq\|(\bm{\Psi}\hat{\bm{T}})_{i,r_{\sigma}+1:d_{0}}\|\ \text{for all}\ \ i\in[r_{\sigma}].$
@@
Let $(\bm{\Psi}\hat{\bm{T}})_{r_{\sigma}+1:d_{L},r_{\sigma}+1:d_{0}}=\bm{P}\bm{\Lambda}\bm{Q}^{T}$ be an SVD of $(\bm{\Psi}\hat{\bm{T}})_{r_{\sigma}+1:d_{L},r_{\sigma}+1:d_{0}}$, where $\bm{\Lambda}\in\mathbb{R}^{(d_{L}-r_{\sigma})\times(d_{0}-r_{\sigma})}$ is diagonal with entries $\gamma_{1}\geq\dots\geq\gamma_{d_{\min}-r_{\sigma}}$ being the top $d_{\min}-r_{\sigma}$ singular values. To proceed, we define
@@
$\bm{\beta}:=\left(\varphi(\sigma_{1}(\bm{W}_{L})),\dots,\varphi(\sigma_{r_{\sigma}}(\bm{W}_{L})),\gamma_{1},\dots,\gamma_{d_{\min}-r_{\sigma}}\right)\in\mathbb{R}^{d_{\min}}.$ (100)
@@
Let $\bm{\Pi}\in\mathcal{P}^{d_{\min}}$ be a permutation matrix such that the entries $\bm{\Pi}\bm{\beta}$ are in a decreasing order and
$\bar{\bm{U}}_{L}=\bm{U}_{L}\mathrm{BlkDiag}\left(\bm{I}_{r_{\sigma}},\bm{P}\right)\mathrm{BlkDiag}\left(\bm{\Pi}^{T},\bm{I}_{d_{L}-d_{\min}}\right),$ (101)
$\bar{\bm{V}}_{1}=\bm{V}_{1}\hat{\bm{T}}\mathrm{BlkDiag}\left(\bm{I}_{r_{\sigma}},\bm{Q}\right)\mathrm{BlkDiag}\left(\bm{\Pi}^{T},\bm{I}_{d_{0}-d_{\min}}\right),$
$\bar{\bm{Y}}=\mathrm{BlkDiag}\left(\bm{\Pi}\mathrm{diag}\left(\bm{\beta}\right)\bm{\Pi}^{T},\bm{0}_{(d_{L}-d_{\min})\times(d_{0}-d_{\min})}\right).$
@@
Since both $\bm{Y}$ and $\bar{\bm{Y}}$ are diagonal matrices with elements in decreasing order, by using Mirsky’s inequality (see Lemma D.8 in Appendix D), we have
[…]
@@
where the equality follows from (140), (100), and (101). This, together with (86), implies
$\|\bm{Y}-\bar{\bm{Y}}\|_{F}\leq\frac{\delta_{y}}{3},\qquad\|\bar{\bm{Y}}\|\leq\|\bar{\bm{Y}}-\bm{Y}\|+\|\bm{Y}\|\leq y_{1}+\frac{\delta_{y}}{3}.$ (103)
Using the triangle inequality, we obtain
$\quad\quad\|\bar{\bm{U}}_{L}^{T}\bm{Y}\bm{Y}^{T}\bar{\bm{U}}_{L}-\bar{\bm{Y}}\bar{\bm{Y}}^{T}\|_{F}$
$~~~\leq\|(\bar{\bm{U}}_{L}^{T}\bm{Y}\bar{\bm{V}}_{1}-\bar{\bm{Y}})\bar{\bm{V}_{1}}^{T}\bm{Y}^{T}\bar{\bm{U}}_{L}\|_{F}+\|\bar{\bm{Y}}(\bar{\bm{U}}_{L}^{T}\bm{Y}\bar{\bm{V}}_{1}-\bar{\bm{Y}})^{T}\|_{F}$
[…]
Next, we write $\bar{\bm{U}}_{L}$ and $\bm{I}_{d_{L}}$ as follows:
$\bar{\bm{U}}_{L}=\begin{bmatrix}(\bar{\bm{U}}_{L}^{(1)})^{T}&\cdots&(\bar{\bm{U}}_{L}^{(p_{Y})})^{T}&(\bar{\bm{U}}_{L}^{(p_{Y}+1)})^{T}\end{bmatrix}^{\!T},\ \bm{I}_{d_{L}}=\begin{bmatrix}(\bm{E}^{(1)})^{T}&\cdots&(\bm{E}^{(p_{Y})})^{T}&(\bm{E}^{(p_{Y}+1)})^{T}\end{bmatrix}^{\!T},$
$\bm{E}^{(1)}=[\,\bm{I}_{h_{1}}\ \ \bm{0}_{h_{1}\times(d_{L}-s_{1})}\,],\ \bm{E}^{(i)}=[\,\bm{0}_{h_{i}\times s_{i-1}}\ \ \bm{I}_{h_{i}}\ \ \bm{0}_{h_{i}\times(d_{L}-s_{i})}\,],\ \bm{E}^{(p_{Y}+1)}=[\,\bm{0}_{(d_{L}-r_{Y})\times r_{Y}}\ \ \bm{I}_{d_{L}-r_{Y}}\,],$
where $\bar{\bm{U}}_{L}^{(i)},\ \bm{E}^{(i)}\in\mathbb{R}^{h_{i}\times d_{L}}$ for all $i\in[p_{Y}]$, and $\bar{\bm{U}}_{L}^{(p_{Y}+1)},\ \bm{E}^{(p_{Y}+1)}\in\mathbb{R}^{(d_{L}-r_{Y})\times d_{L}}$. Applying the Davis-Kahan theorem (see Lemma D.7 in Appendix D) to the matrices $\bar{\bm{U}}_{L}^{T}\bm{Y}\bm{Y}^{T}\bar{\bm{U}}_{L}$ (considered as $\bm{A}$), $\bar{\bm{Y}}\bar{\bm{Y}}^{T}$ (considered as $\hat{\bm{A}}$), $\bar{\bm{U}}_{L}^{(i)^{T}}$ (considered as $\bm{V}$) , and $\bm{E}^{(i)^{T}}$ (considered as $\hat{\bm{V}}$) for each $i\in[p_{Y}+1]$, there exist orthogonal matrices $\hat{\bm{U}}_{L}^{(i)}\in\mathcal{O}^{h_{i}}$ for each $i\in[p_{Y}]$ and $\hat{\bm{U}}_{L}^{(p_{Y}+1)}\in\mathcal{O}^{d_{L}-r_{Y}}$ such that
where $\lambda_{i}$ is the $i$-th largest non-repeated eigenvalue of $\bar{\bm{U}}_{L}^{T}\bm{Y}\bm{Y}^{T}\bar{\bm{U}}_{L}$ and $\lambda_{0}=\infty$ and $\lambda_{p_{Y}+2}=$-$\infty$ by convention. We compute $\lambda_{i}=y_{s_{i}}^{2}$ for $i\in[p_{Y}]$ and $\lambda_{p_{Y}+1}=0$ for $i\in[p_{Y}]$. Using (104) and the following inequalities